A sofa and a chair cost $850 as a set. The sofa cost $100 more than twice as much as the chai. Write 2 linear equations and what is the cost of the sofa and the cost of the chair.
Answer by stanbon(75887)

A sofa and a chair cost $850 as a set. The sofa cost $100 more than twice as much as the chai. Write 2 linear equations and what is the cost of the sofa and the cost of the chair.
----------------------------------------
EQUATIONS:
s + c = 850
s = 2c + 100
----------------
Rearrange:
s = -c + 850
s = 2x+100
----------------
Solve for "c".
-c+850 = 2c+100
3c = 750
c = $250 (cost of a chair)
------------
s = 250 = 850
s = $600 (cost of a sofa)
=====================
Cheers,
Stan H.